OpenStudy (anonymous):

Use The Inequality -4t-8≤12 To Fill In The Missing Numbers . A.t≥? B.t+4≥? C.t-?≥0 D.t+10≥? E.3t≥? F.t≥-5 -

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

that was wrong, lets do it right

OpenStudy (anonymous):

Lol Okay

OpenStudy (anonymous):

\[-4t-8≤12 \] add 8 get \[-4t\leq 20\] divide by -4 and change the inequality get \[t\geq -5\]

OpenStudy (anonymous):

then if \[t\geq -5\] adding 4 to both sides gives \[t+4\geq -1\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So What About C,D,E,F

OpenStudy (anonymous):

if you want 0 on the right, add 5 to both sides to get \[t+5\geq 0\]

OpenStudy (anonymous):

since \[t\geq -5\] add 15 to get \[t+15\geq 10\]